OpenStudy (anonymous):

5.6% is what as a fraction

OpenStudy (anonymous):

\[\frac{ 56 }{ 100 } simplified \frac{ 23 }{ 50 }\]

OpenStudy (anonymous):

5.6x1/100=0.056 answer=0-056. this is correct or not. i am not confirm.